If I buy a square lot that is 110 yards on all four sides, what is the size in acres?

I bought a pound of nails. If each nail weighs 0.025 ounces how many nails are there total?

Answer by stanbon(75887)

If I buy a square lot that is 110 yards on all four sides, what is the size in acres?
Area= s^2 = 110^2 = 12100 sq yrds.















Area = 110^2 = 12100 sq. yrds
--------------------------------
16 ounces / 0.025 ounces = 640 nails
Cheers,
Stan H.

















































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































































If I buy a square lot that is 110 yards on all four sides, what is the size in acres?
Area=110^2 = 12100 sq yrds.
I bought a pound of nails. If each nail weighs 0.025 ounces how many nails are there total?
16 ounces /0.025 = 40 nails.
Cheers,
Stan H.



















I bought a pound of nails. If each nail weighs 0.025 ounces how many nails are there total?
1 lb / 0.025 ounces = 16 ounces/ 0.025 ounces = 640 nails.
Cheers,
Stan H.